OpenStudy (anonymous):

WILL MEDAL The sides of a square are three to the power of two sevenths inches long. What is the area of the square? nine to the power of four sevenths square inches three to the power of four sevenths square inches three to the power of the fraction four over forty nine square inches nine to the power of the fraction four over forty nine square inches

OpenStudy (anonymous):

@sweetburger

OpenStudy (danjs):

\[\large a^b*a^c = a^{b+c}\]

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (danjs):

or \[\large [a^b]^c = a^{b*c}\]

OpenStudy (anonymous):

so its b

OpenStudy (danjs):

yeah, area is the length squared

OpenStudy (anonymous):

ok so im right its the second choice

OpenStudy (danjs):

yes \[\large [3^{2/7}]^2 = 3^{2 * 2/7} = 3^{4/7}\]

OpenStudy (anonymous):

thx